OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary round to the nearest hundredth. x^2-6=x

OpenStudy (anonymous):

Let's solve your equation step-by-step. x2-6=x Step 1: Subtract x from both sides. x2-6-x=x-x x2-x-6=0 Step 2: Factor left side of equation. (x+2)(x-3)=0 Step 3: Set factors equal to 0. x+2=0 or x-3=0 x=-2 or x=3

OpenStudy (anonymous):

Thank you SO much!!! :D

OpenStudy (anonymous):

yupp